# From Priest to Doctor: Domain Adaptation for Low-Resource Neural Machine Translation

### v1

@@ Abstract
Ali Marashian1, Enora Rice1, Luke Gessler2, Alexis Palmer1, Katharina von der Wense1,3 1University of Colorado Boulder, 2Indiana University Bloomington, 3Johannes Gutenberg University Mainz Correspondence: ali.marashian@colorado.edu
@@ 1 Introduction
Neural machine translation (NMT) models have limited ability to deal with languages that lack large-scale monolingual and parallel corpora Wang et al. (2021). Moreover, NMT systems face challenges when translating text from a novel domains characterized by unique style or vocabulary Koehn and Knowles (2017); Saunders (2022). Often, these issues co-occur, a scenario that has been neglected by researchers so far. Most of the world’s 7000+ languages are considered low-resource Joshi et al. (2020), and existing data for them are in limited domains; the languages that could most benefit from domain adaptation (DA) are the ones left behind.
Figure 1: The (neglected) intersection of Low-Resource NMT and DA in NMT, has led us to consider only these commonly accessible resources.
In this paper, we explore a realistic setting in which we aim to translate between a high-resource and a low-resource language and are restricted to the following commonly available resources: a) Bible translation, i.e., a small parallel corpus in the source domain; b) monolingual target-domain texts in the high-resource language; and c) a bilingual dictionary for the two languages. To keep the setting generalizable, we assume neither access to a model pretrained on text in the low-resource language nor access to data in a related high-resource language, as for many truly low-resource languages, those are impossible to find.
We experiment with a set of four DA and low-resource NMT methods and aim to translate from English to a simulated low-resource target language. We use mBART Liu et al. (2020) which has been finetuned on parallel Bible texts as our base model, and our goal is to adapt to the target domains of government documents and medicine. The methods we investigate use the bilingual dictionaries in various ways.
Our experiments showcase the varying effectiveness of existing methods: the weakest approach results in models that perform worse than the base model, while the best approach – which, surprisingly, is also the simplest – results in a ChrF score more than twice as high as the base model’s. However, as the best model only reaches a ChrF score of 42.47 and a BLEU score of 13.47 (on average), we also perform a small human evaluation, which confirms that there is still a need for the development of better DA methods for low-resource NMT.
@@ 2 Related Work
As domains are defined by the characteristics of data Saunders (2022), many effective DA approaches focus on the data and thus can be applied to various underlying architectures. Some works focus on acquiring monolingual in-domain data, which is easier to find than in-domain parallel data. Back-translation uses monolingual target-domain data in the target language and produces artificial source sentences using a target-to-source NMT model Poncelas et al. (2019); Jin et al. (2020). Chinea-Ríos et al. (2017) use monolingual source-side corpora and a source-to-target NMT model for forward-translation, where it is common to employ self-learning. With access to a small parallel corpus, extra training data can be created by introducing noise Vaibhav et al. (2019). Synthetic parallel data can be acquired from an external source or generated using a predefined or induced lexicon. Hu et al. (2019) use a lexicon to back-translate target-side sentences. Peng et al. (2020) use a dictionary, injecting dictionary terms into out-of-domain texts to synthesize in-domain training data. Bergmanis and Pinnis (2021) augment the training data by annotating randomly selected source language words with their target language lemmas to integrate terms. Zhang et al. (2022) introduce lexical constraints into iterative back-translation.
@@ 3 Data
In all experiments, the only parallel data we use for training is the JHU Bible Corpus McCarthy et al. (2020).
@@
We explore adapting to two different domains, one at a time: government documents and medicine. The domain-specific data mostly come from past WMT translation tasks Barrault et al. (2020); Akhbardeh et al. (2021); Kocmi et al. (2022, 2023). As we only assume monolingual in-domain training data (cf. Section 4), training and pretraining use only source-side sentences from these parallel data sets. Data availability varies across language/domain pairs, and we cap data set sizes to maintain comparability across languages. For training we use no more than 200K sentence pairs. If our setting requires pretraining, we use the same source-side sentences used for training. For testing, we use 1500 sentence pairs.
@@ 5 Results and Discussion
CPT is helpful in most of the experiments when compared with plain mBART, but not compared to DALI. After pretraining, the model is fine-tuned on only the biblical data. In the pretraining, we reconstruct the source side, so the model only learns to output in the target language from the Bible verses. Pretraining helps the model get more familiar with the domain and establish connections between the embeddings of target words and their respective translations in the high-resource language. [^4] According to Liu et al. (2021), the performance boost is expected to increase if we have monolingual texts in the target language instead and can use them during pretraining.
Combining the methods together shows some improvements over other individual methods, but generally fails to reach DALI’s performance. Note that the same dataset was used to both pretrain the model (the CPT part), and then make pseudo-parallel data (the DALI part). Since LeCA is not helpful when added to the basic mBART, we also test performance of Combined without LeCA, on the medical domain. The results (Table 7) indicate that - when using the same dataset for both - adding pretraining on top of DALI can be detrimental, but removing LeCA increases performance on all languages for the medical domain.
@@
We see some interesting trends in the outputs. For the Maltese-medical test set, the language-domain pair with the most significant performance boost, both mBART and LeCA translate in a religious tone (see Table 5 in App. C for one source sentence and multiple outputs). The same is true for CPT, which also tends to copy words from the input - as it was a part of the reconstruction procedure during pretraining. It is important to emphasize that Maltese is a morphologically rich language, and the inflections are mostly discarded in the outputs of DALI and Combined; for example the words are more likely to be disjoint in their outputs than they are in the target (the first “il $\hbar$ra” vs “$\hbar$ra” in the target), or they can be in different forms (“huwa” vs “hija”). Maltese at times has a different word order than that of English ( “il-mediċina $\hbar$ajnejn” is translated as “$\hbar$ajn mediċina”, which matches the order of its English counterpart “eye medicine”), and it is also more flexible. DALI and Combined produce word orders that closely follow the source language.
@@ 6 Conclusion
As our research shows that these methods do not sufficiently enhance performance of the models for the models to be deemed useful, there are some caveats to be mindful of. Specifically, these methods should not be used for real-world MT in critical contexts involving low-resource languages; e.g. providing medical advice based on the translations produced by the model.
@@ Appendix B Training
We will publish the code base upon the acceptance of the paper.
@@ Appendix C Additional Outputs and Evaluation
Source: if the other eye medicine is an eye ointment it should be used last .
mBART: jekk il-mara $\hbar$or hi çajn $\hbar$ra , hi çandha tinçatalha $\hbar$$\hbar$ar $\hbar$$\hbar$ar . LeCA: inkella jekk il-mara $\hbar$or hi żejt , tkun maçmula $\hbar$$\hbar$ar .
DALI: jekk il $\hbar$ra $\hbar$ajn mediċina huwa an $\hbar$ajn infużjoni dan $\hbar$andu tkun użati $\hbar$$\hbar$ar . CPT: jekk $\hbar$rajn ta ’ $\hbar$or hi çajnejja ointment , it should be used $\hbar$$\hbar$ar .
Combined: jekk il $\hbar$ra $\hbar$ajn mediċina huwa an $\hbar$ajn ointment dan $\hbar$andu tkun użati $\hbar$$\hbar$ar . Target: jekk il-mediċina $\hbar$ajnejn $\hbar$ra hija ingwent $\hbar$ajnejn , dan $\hbar$andu jintuża $\hbar$$\hbar$ar .
Table 5: An example of different model outputs for a Maltese sentence in the medical domain.
Table 5 is discussed in Section 4.

### v3

@@ Abstract
Ali Marashian,1 Enora Rice,1 Luke Gessler,2 Alexis Palmer,1 Katharina von der Wense1,3 1University of Colorado Boulder, 2Indiana University Bloomington, 3Johannes Gutenberg University Mainz ali.marashian@colorado.edu
@@ 1 Introduction
Neural machine translation (NMT) models have limited ability to deal with languages that lack large-scale monolingual and parallel corpora Wang et al. (2021). Moreover, NMT systems face challenges when translating text from novel domains characterized by unique style or vocabulary Koehn and Knowles (2017); Saunders (2022). Often, these issues co-occur, a scenario that has been neglected by researchers so far. Most of the world’s 7000+ languages are considered low-resource Joshi et al. (2020), and existing data for them are in limited domains; the languages that could most benefit from domain adaptation (DA) are the ones left behind.
Figure 1: In our work, which looks at the (previously neglected) intersection of low-Resource NMT and domain adaptation in NMT, we consider only these commonly accessible resources.
In this paper, we explore a realistic setting in which we aim to translate between a high-resource and a low-resource language and are restricted to the following commonly available resources: a) Bible translations, i.e., a small parallel corpus in the source domain; b) monolingual target-domain texts in the high-resource language; and c) a bilingual dictionary for the two languages. To keep the setting generalizable, we assume neither access to a model pretrained on text in the low-resource language nor access to data in a related high-resource language, as for many truly low-resource languages, those are impossible to find.
We experiment with a set of four DA and low-resource NMT methods and aim to translate from English to a target language, simulating a low-resource setting. We use mBART Liu et al. (2020) which has been fine-tuned on parallel Bible texts as our base model, and our goal is to adapt it to the target domains of government documents and medicine. The methods we investigate use the bilingual dictionaries in various ways.
Our experiments showcase the varying effectiveness of existing methods: the weakest approach results in models that perform worse than the base model, while the best approach – which, surprisingly, is also the simplest – results in a ChrF score more than twice as high as the base model’s. However, as the best model only reaches a ChrF score of 42.47 and a BLEU score of 13.47 (on average), we also perform a small human evaluation, which confirms that there is still a need for the development of better DA methods for low-resource NMT. Our code is available on GitHub. [^1] https://github.com/alimrsn79/da_lr_nmt
@@ 2 Related Work
As domains are defined by the characteristics of data Saunders (2022), many effective DA approaches focus on the data and, thus, can be applied to various underlying architectures. Some works focus on acquiring monolingual in-domain data, which is easier to find than in-domain parallel data. Back-translation uses monolingual target-domain data in the target language and produces artificial source sentences using a target-to-source NMT model Poncelas et al. (2019); Jin et al. (2020). Chinea-Ríos et al. (2017) use monolingual source-side corpora and a source-to-target NMT model for forward-translation, where it is common to employ self-learning. With access to a small parallel corpus, extra training data can be created by introducing noise Vaibhav et al. (2019). Synthetic parallel data can be acquired from an external source or generated using a predefined or induced lexicon. Hu et al. (2019) use a lexicon to back-translate target-side sentences. Peng et al. (2020) use a dictionary, injecting dictionary terms into out-of-domain texts to synthesize in-domain training data. Bergmanis and Pinnis (2021) augment the training data by annotating randomly selected source language words with their target language lemmas to integrate terms. Zhang et al. (2022) introduce lexical constraints into iterative back-translation.
@@ 3 Data
In all experiments, the only parallel data we use for training come from the JHU Bible Corpus McCarthy et al. (2020).
@@
We explore adapting to two different domains, one at a time: government documents and medicine. The domain-specific data mostly come from past WMT translation tasks Barrault et al. (2020); Akhbardeh et al. (2021); Kocmi et al. (2022, 2023). As we assume only monolingual in-domain training data (cf. Section 4), training and pretraining use only source-side sentences from these parallel data sets. Data availability varies across language/domain pairs, and we cap data set sizes to maintain comparability across languages. For training we use no more than 200K sentence pairs. If our setting requires pretraining, we use the same source-side sentences used for training. For testing, we use 1500 sentence pairs.
@@ 5 Results and Discussion
CPT is helpful in most of the experiments when compared to plain mBART, but not compared to DALI. After pretraining, the model is fine-tuned only on Bible data. In the pretraining, we reconstruct the source side, so the model only learns to output in the target language from the Bible verses. Pretraining helps the model get more familiar with the domain and establish connections between the embeddings of target words and their respective translations in the high-resource language. [^5] According to Liu et al. (2021), the performance boost is expected to increase if we have monolingual texts in the target language instead and can use them during pretraining.
Combining the methods together shows some improvements over other individual methods, but generally fails to reach DALI’s performance. Note that the same dataset was used to both pretrain the model (the CPT part) and to then make pseudo-parallel data (the DALI part). Since LeCA is not helpful when added to the basic mBART, we also test performance of Combined without LeCA, on the medical domain. The results (Table 8) indicate that – when using the same dataset for both – adding pretraining on top of DALI can be detrimental, but removing LeCA increases performance on all languages for the medical domain.
LeCA only uses the dictionary in the final stage, and CPT uses the monolingual data during pretraining, before being fine-tuned on the bible data. DALI and Combined are the only methods that have access to source-side target-domain monolingual data during the final stage of training, which could partially explain their superior performance. That a simple method like DALI – that mostly keeps the word order of the sentence language – should be the best performing method hints at the extensive room for growth in future work.
Figure 2 shows averaged sentence-level BLEU and ChrF scores plotted against their respective reference token lengths for DALI models. For length $l$, we average the scores of the models for different languages if the reference translation is of length $l$. We can see that generally the scores seem to get higher with longer sentences, especially for ChrF.
Figure 2: The trend of averaged sentence-level BLEU (left) and ChrF (right) scores against the token length of the reference translation for DALI models. The scores are averaged across all the model outputs of the same length – including averaging across languages, where relevant.
@@
We see some interesting trends in the outputs. Table 2 showcases an example with the outputs of different methods for one sentence from the Maltese-medical test set, the language–domain pair with the most significant performance boost. Warning: these outputs could include distressing language against women that may harm some readers. Both mBART and LeCA translate in a religious tone. The same is true for CPT, which also tends to copy words from the input – as it was a part of the reconstruction procedure during pretraining. It is important to emphasize that Maltese is a morphologically rich language, and the inflections are mostly discarded in the outputs of DALI and Combined; for example the words are more likely to be disjoint in their outputs than they are in the target (the first “il $\hbar$ra” vs “$\hbar$ra” in the target), or they can be in different forms (“huwa” vs “hija”). Note that ointment was translated to infusion by DALI. Given the sensitivity of the domain, a translation like this can potentially be harmful.
Source: if the other eye medicine is an eye ointment it should be used last
mBART: jekk il-mara $\hbar$or hi çajn $\hbar$ra , hi çandha tinçatalha $\hbar$$\hbar$ar $\hbar$$\hbar$ar LeCA: inkella jekk il-mara $\hbar$or hi żejt , tkun maçmula $\hbar$$\hbar$ar
BT: if the other woman is someone else, she should be punished in the end BT: otherwise if the other woman is a virgin, she will be the worst
DALI: jekk il $\hbar$ra $\hbar$ajn mediċina huwa an $\hbar$ajn infużjoni dan $\hbar$andu tkun użati $\hbar$$\hbar$ar CPT: jekk $\hbar$rajn ta ’ $\hbar$or hi çajnejja ointment , it should be used $\hbar$$\hbar$ar
BT: If the other eye medicine is an eye infusion, this should be used last BT: Even though the other one is a çajnejja ointment, it should be used last
Combined: jekk il $\hbar$ra $\hbar$ajn mediċina huwa an $\hbar$ajn ointment dan $\hbar$andu tkun użati $\hbar$$\hbar$ar Target: jekk il-mediċina $\hbar$ajnejn $\hbar$ra hija ingwent $\hbar$ajnejn , dan $\hbar$andu jintuża $\hbar$$\hbar$ar
BT: If the other eye medicine is an eye ointment, this should be used last BT: Although the other eye medicine is an eye ointment, this should be used last
Table 2: Warning: this table contains harmful language about women that may distress some readers. An example of different model outputs for a Maltese sentence in the medical domain. For better comparison, the back-translations (BT) of the outputs to English are also included, done via Google Translate.
Maltese at times has a different word order than that of English ( “il-mediċina $\hbar$ajnejn” is translated as “$\hbar$ajn mediċina”, which matches the order of its English counterpart “eye medicine”), and it is also more flexible. DALI and Combined produce word orders that closely follow the source language.
@@ 6 Conclusion
As our research shows that these methods do not sufficiently enhance performance for the models to be deemed useful, there are some caveats to be mindful of. Specifically, these methods should not be used for real-world MT in critical contexts involving low-resource languages; e.g. providing medical advice based on the translations produced by the model.
@@
Acknowledgments
We sincerely thank the anonymous reviewers for their helpful feedback. We also wish to express our gratitude to LECS and NALA labs at the University of Colorado Boulder. This work made use of the Blanca condo computing resource at the University of Colorado Boulder, which is supported through joint funding by computing users and the university.
@@ Appendix C Additional Outputs and Evaluation
Output statistics
The average number of words, number of tokens, and number of characters of the outputs of different methods against the reference translations are presented in Table 7. For number of words, an output is split by white-spaces. For tokens, the mBART tokenizer is used. We average the results across languages. We report the averages because relative length patterns tend to be consistent across languages. The full table containing language specific statistic is available on the GitHub repository: https://github.com/alimrsn79/da_lr_nmt.
@@
Average
Domain Words Tokens Characters
Reference Gov. 23.51 44.65 154.55
Med. 19.65 39.43 120.42
mBART Gov. 25.12 49.25 132.7
Med. 20.91 44.27 110.94
DALI Gov. 26.83 43.99 155.28
Med. 20.3 37.38 115.7
LeCA Gov. 24.82 49.15 133.92
Med. 22.94 46.02 117.94
CPT Gov. 26.44 47.85 139.7
Med. 20.5 38.23 109.46
Combined Gov. 26.93 43.97 155.95
Med. 20.31 37.41 116.09
Table 7: The average number of words, tokens, and characters of the outputs of different methods against the reference translation. The results are averaged over all the experiments.